(4) Evaluate the integral. (Hint: Substitution Rule] |(2 (2x + 3)(2x2 + 6x + 1)8dx
Answer #1

Solution: - Let I I= f(2012) (2012+62+1)8221 Let 2012+62+1 t (4x+6) dx = dt (22+3) dx = - dut I = s Lujo (tadt) {by by substi

Thanks and Happy learning.

ask me if you have any doubts.......please rate my answer....